OpenStudy (anonymous):

-3 + 4 = 5x - 6. Is the answer -1?

OpenStudy (studygurl14):

-3 + 4 = 5x - 6 1 = 5x - 6 7 = 5x You tell me. Is -1 correct, @Timg ?

OpenStudy (anonymous):

Sorry it should have been -3x +4 = 5x - 6

OpenStudy (studygurl14):

I'm confused about what you mean @Timg

OpenStudy (anonymous):

I forgot to add the x on the 3 in my first question.

OpenStudy (bmorse):

8x=10 x=5/4

OpenStudy (studygurl14):

Where'd you get that, @bmorse ?

OpenStudy (anonymous):

Ahh. Add 3x to 5x and add 6 to the 4. Then you get the 8x=10. Right? Got it. Thanks.

OpenStudy (studygurl14):

Oh, I get it. You forgot to put the 3 on the x, which changes the equation completely.

OpenStudy (anonymous):

Sorry about that. Can't do math or type !

OpenStudy (bmorse):

yup!